# CP-uniGuard: A Unified, Probability-Agnostic, and Adaptive Framework for Malicious Agent Detection and Defense in Multi-Agent Embodied Perception Systems

### v2

@@ Abstract
Collaborative Perception (CP) has been shown to be a promising technique for multi-agent autonomous driving and multi-agent robotic systems, where multiple agents share their perception information to enhance the overall perception performance and expand the perception range. However, in CP, an ego agent needs to receive messages from its collaborators, which makes it vulnerable to attacks from malicious agents. To address this critical issue, we propose a unified, probability-agnostic, and adaptive framework, namely, CP-uniGuard, which is a tailored defense mechanism for CP deployed by each agent to accurately detect and eliminate malicious agents in its collaboration network. Our key idea is to enable CP to reach a consensus rather than a conflict against an ego agent’s perception results. Based on this idea, we first develop a probability-agnostic sample consensus (PASAC) method to effectively sample a subset of the collaborators and verify the consensus without prior probabilities of malicious agents. Furthermore, we define collaborative consistency loss (CCLoss) for object detection task and bird’s eye view (BEV) segmentation task to capture the discrepancy between an ego agent and its collaborators, which is used as a verification criterion for consensus. In addition, we propose online adaptive threshold via dual sliding windows to dynamically adjust the threshold for consensus verification and ensure the reliability of the systems in dynamic environments. Finally, we conduct extensive experiments and demonstrate the effectiveness of our framework. Code will be released at https://github.com/CP-Security/CP-uniGuard.
@@ 1 Introduction
The remainder of this paper is organized as follows. In Section 2, we review the related works on collaborative perception, adversarial perception, and defensive perception. In Section 3, we provide the formulation and pipeline of collaborative perception, and then we discuss the adversarial threat model. In Section 4, we propose our unified, probability-agnostic, and adaptive framework, CP-uniGuard, which consists of a probability-agnostic sample consensus (PASAC) method and two collaborative consistency losses (CCLoss) for the object detection task and BEV segmentation task. In addition, we also propose an online adaptive threshold scheme to dynamically adjust the threshold for consensus verification, which can ensure the reliability of the system in dynamic environments. In Section 5, we conduct extensive experiments to evaluate the performance of our framework. Finally, we discuss the limitations and future work in Section 6 and conclude this paper in Section 7.
2 Background and Related Work
2.1 Collaborative Perception
@@ 3 Problem Setup
6:for $i=0$ to $n-1$ do
7:     $\mathbf{f}_{i,t}\leftarrow E_{\theta_{i}}(\mathbf{x}_{i,t})$triangleright Extract features
@@
10:$\mathtt{Mess}_{t}\leftarrow\emptyset$triangleright Initialize message set for ego agent
11:for $i=1$ to $n-1$ do
12:     $\mathbf{f}_{i,t}^{w}\leftarrow\mathcal{W}(\mathbf{f}_{i,t},\mathbf{T}_{i\to 0,t})$triangleright Warp features
13:     $\mathbf{m}_{i,t}\leftarrow C_{\phi}(\mathbf{f}_{i,t}^{w})$triangleright Compress features (optional)
14:     $\mathtt{Mess}_{t}\leftarrow\mathtt{Mess}_{t}\cup\{\mathbf{m}_{i,t}\}$triangleright Add to message set
@@
17:$\mathbf{z}_{t}\leftarrow A_{\psi}(\mathbf{f}_{0,t},\mathtt{Mess}_{t})$triangleright Fuse the ego’s and collaborators’ features
@@
19:$\hat{\mathbf{y}}_{t}\leftarrow D_{\omega}(\mathbf{z}_{t})$triangleright Decode the fused features into the perception output
@@
3.1 Collaborative Perception
As stated in Section 2.1, CP can be categorized into three classes according to the stage of collaboration, including early fusion, intermediate fusion, and late fusion. In this section, we focus on the widely used intermediate fusion paradigm, which is the most efficient and effective fusion strategy.
@@ 3.1 Collaborative Perception
3.2 Adversarial Threat Model in Collaborative Perception
@@
where $\boldsymbol{\delta}_{i,t}$ is the adversarial perturbation. $\mathcal{B}_{p}(\Delta)=\{x:\lVert x\rVert_{p}\leq\Delta\}$ is an $\ell_{p}$ ball to constrain the magnitude of the perturbation. Unless stated otherwise, we adopt the imperceptibility norm $p=\infty$ with $\Delta=0.1$
@@ 4 CP-uniGuard Framework
14:     $\hat{\mathbf{y}}_{\mathrm{G}_{1},t}\leftarrow D_{\omega}(A_{\psi}(\mathbf{f}_{0,t},\{\mathbf{f}_{i,t}^{w}\}_{i\in\mathrm{G}_{1}}))$triangleright Fusion
15:     $\hat{\mathbf{y}}_{\mathrm{G}_{2},t}\leftarrow D_{\omega}(A_{\psi}(\mathbf{f}_{0,t},\{\mathbf{f}_{i,t}^{w}\}_{i\in\mathrm{G}_{2}}))$triangleright Fusion
@@
19:     $\triangleright$ First group contains malicious agents
@@
25:     $\triangleright$ Second group contains malicious agents
@@
4.1 PASAC: Probability-Agnostic Sample Consensus
4.1.1 Our Method
@@
More specifically, the ego agent will generate the collaborative perception results $\mathbf{y}_{t}$ based on its observation and the received messages for feature fusion from the $n-1$ other agents. Firstly, the ego agent generates its perception results $\mathbf{y}_{0,t}$ based on its own observation. Then, it randomly splits the agents into two groups $\mathrm{G}_{1}$ and $\mathrm{G}_{2}$ of equal size. After receiving all the messages, the ego agent fuses the features and generates the perception results $\mathbf{y}_{\mathrm{G}_{1},t}$ based on the messages from the first group and $\mathbf{y}_{\mathrm{G}_{2},t}$ based on the messages from the second group.
@@ 4.1.2 Theoretical Analysis
In the worst-case scenario, we disregard the early stopping condition $k=n-m$, which is met when all benign agents are identified. We define the positive dyadic-logarithm $\bigl{\lceil}\log_{2}x\bigr{\rceil}_{+}$ as $\max(0,\lceil\log_{2}x\rceil)$.
@@
[…]
Every benign singleton $\{i\}$ with $i\notin\mathcal{M}$ is queried exactly once and produces a benign leaf. Hence there are precisely $n-m$ such leaves. Adding the two contributions yields
@@
4.2 Collaborative Consistency Loss Verification
@@
4.2.1 Collaborative Consistency Loss for Object Detection
@@
4.2.2 Collaborative Consistency Loss for BEV Segmentation
@@
4.3 Online Adaptive Threshold via Dual Sliding Windows
@@
For any window $\mathcal{W}$ of size $n$, denote the empirical $q$‑quantile by $\hat{q}_{q}(\mathcal{W})=\inf\bigl{\{}z:\tfrac{1}{n}\sum_{Z\in\mathcal{W}}\!\mathbf{1}\{Z\leq z\}\geq q\bigr{\}}.$ We use the upper $(1-\alpha)$‑quantile of benign scores and the lower $\beta$‑quantile of contaminated scores:
@@ 5 Experiments
TABLE I: Quantitative results on BEV segmentation task. Upper-bound denotes collaborative perception with all benign agents. Lower-bound is individual perception.
@@
5.1 Experimental Setup
5.1.1 Datasets and Evaluation Metrics.
In our experiments, we leverage V2X-Sim [33] as our dataset. It is the first synthetic dataset for CP generated by CARLA-SUMO co-simulator. In addition, to evaluate the performance of the object detection and BEV segmentation tasks, we adopt average precision (AP) with Intersection over Union (IoU) thresholds 0.5 and 0.7 (AP@0.5 and AP@0.7) for object detection task, and mean Intersection over Union (mIoU) for BEV segmentation task. We also use Verification Count to evaluate the performance of PASAC, which is the total number of times that malicious agents are checked.
5.1.2 Implementation Details.
We leverage the backbone which is the same as [34] for the object detection task, and use U-Net [35] as the backbone for the BEV segmentation task. The fusion method for the both tasks is V2VNet [36]. Our experiment is deployed on a computer consisting of 2 Intel(R) Xeon(R) Silver 4410Y CPUs (2.0GHz), four NVIDIA RTX A5000 GPUs, and 512GB DDR4 RAM. As for the implementation of adversarial attacks, we employ three kinds of attacks: fast gradient sign method (FGSM) [37], Carlini & Wagner (C&W) [38], and the projected gradient descent (PGD) [39]. For each attack, we set the maximum perturbation $\delta_{\mathrm{max}}=0.1$, iterations steps $T=15$, and the step size $\gamma=0.01$.
@@ 5.1 Experimental Setup
5.2 Evaluation and Analysis
5.2.1 Evaluation of CP-uniGuard.
We evaluate the efficacy of our CP-uniGuard scheme against a variety of adversarial attacks. The outcomes of these evaluations are detailed in Table I and Table II. As shown in Table I, in scenarios where the CP system lacks defensive mechanism, the mIoU across all three attack modalities significantly falls below the established lower bound, registering at 37.09. This substantial degradation in performance underscores the effectiveness of the adversarial attacks implemented. Conversely, our CP-uniGuard framework can effectively defend against these attacks and achieve an mIoU that closely approaches the upper bound of 40.45.
Additionally, we evaluate our approach on the object detection task. Table II presents the Average Precision (AP) at thresholds 0.5 and 0.7 for different defense strategies. Our observations indicate that CP-uniGuard consistently achieves the highest AP@0.5 and AP@0.7 across all attack scenarios, demonstrating its superior performance. Specifically, when defending against the PGD attack, CP-uniGuard achieves an AP@0.5 of 80.4 and an AP@0.7 of 78.3, which is higher than ROBOSAC (77.9 and 75.6) by 2.5 point and 2.7 point, respectively. For defending against the C&W attack, CP-uniGuard performs better, achieving an AP@0.5 of 80.2 and an AP@0.7 of 77.6, which is higher than ROBOSAC (74.5 and 71.1) by 5.7 point and 6.5 point, respectively.
These quantitative results confirm that CP-uniGuard not only defends BEV segmentation (mIoU) but also generalizes effectively to object detection, restoring performance close to the all-benign scenario and substantially outperforming both adversarially trained baselines and prior-art methods such as ROBOSAC.
TABLE II: Quantitative results on object detection task. The upper-bound represents collaborative perception with all benign agents, while the lower-bound corresponds to individual perception.
@@ 5.2 Evaluation and Analysis
5.2.2 Evaluation of PASAC.
@@
TABLE III: Comparison results between ROBOSAC and PASAC. In this experiment, the attack ratios are known to the ROBOSAC method.
@@
5.2.3 Comparison Results in Sampling Efficiency.
In order to evaluate the sampling efficiency of PASAC method. Firstly, we conduct a comparsion experiment with the previous state-of-the-art method, ROBOSAC [5]. Following the experiment setup from ROBOSAC, we evaluate the performance of ROBOSAC and PASAC under different attack ratios. In this setting, the attack ratios are known to the ROBOSAC method. The results are shown in Table III. We observe that PASAC outperforms ROBOSAC in terms of the verification count. Specifically, PASAC achieves a lower verification count than ROBOSAC under different attack ratios. For example, when the attack ratio is 0.6, the average verification count of PASAC is 7.59, which is lower than the average verification count of ROBOSAC (8.29), and when the attack ratio is 0.4, the average verification count of PASAC is 6.60, which is much lower than the average verification count of ROBOSAC (10.36). In addition, the results of PASAC are more stable than ROBOSAC, for example, the maximum verification count of PASAC is 8, while the maximum verification count of ROBOSAC is 46, which is much higher than the average verification count.
Secondly, we conduct a comparison experiment between PASAC, ROBOSAC, and Linear Sampling without the prior probabilities of malicious agents. Here linear sampling is a simple sampling method that verifying the collaborators one by one. The results are shown in Fig. 4. Since the attack ratios are unknown to the ROBOSAC method, it need to estimate the attack ratio first, which leads to a sharp decrease in sampling efficiency. From the figure, we can observe that the verification count of ROBOSAC is much higher than PASAC and Linear Sampling, and it increases sharply when the number of malicious agents increases. On the contrary, our PASAC method does not need to know the attack ratio, and its verification count is much lower than ROBOSAC and Linear Sampling in most cases. These results shows the superiority of our PASAC method in sampling efficiency.
@@
Figure 5: Visualization of no defense and defensive CP-uniGuard results on V2X-Sim datasets.
Figure 6: Visualization of the online adaptive threshold mechanism on V2X-Sim datasets.
5.2.4 Study on the Online Adaptive Threshold Mechanism.
@@
In this section, we conducted an analysis on the online adaptive threshold mechanism. As shown in Fig. 6, we plot the adaptive threshold value with different initial thresholds. We can observe that the threshold can quickly converge to a stable value, which is around 0.4. In addition, the threshold is not fixed, but dynamically adjusted based on the current situation to ensure the robustness of the system and a low false positive rate.
@@
TABLE IV: Generalization Evaluation. We evaluate and compare the generalization capability of CP-uniGuard on two attacks, PGD and C&W, respectively.
@@
5.2.5 Generalization Evaluation
We further evaluate the generalization capability of CP-uniGuard across different attack types, comparing it with feature-level defense methods such as CP-Guard+ [8]. For CP-Guard+, we train the model on the CP-GuardBench dataset [8] while excluding specific attacks from the training set to assess its generalization performance. Specifically, we exclude the PGD attack and C&W attack during training and subsequently evaluate the model’s performance against these two attacks, respectively. Results are presented in Table IV, we can see that CP-uniGuard consistently achieves superior AP@0.5 and AP@0.7 scores compared to CP-Guard+ under both PGD and C&W attacks which are unseen. This performance gap stems from CP-uniGuard’s fundamental design: rather than relying on training, CP-uniGuard leverages consensus verification, which provides inherent generalization to unseen attacks without requiring prior exposure during training.
5.2.6 Qualitative Evaluation
As depicted in Fig. 5, we present the visualization results on the V2X-Sim dataset. Without CP-uniGuard, attackers can significantly disrupt collaborative perception, leading to a marked degradation in the performance of BEV segmentation tasks. However, our introduced CP-uniGuard framework can intelligently identify benign collaborators and eliminate malicious collaborators, thereby facilitating robust CP.
6 Limitations and Future Work
Despite the promising results of CP-uniGuard in defending against malicious agents in collaborative perception, the computational complexity of PASAC, though more efficient than ROBOSAC and linear sampling which checking every agent individually, still introduces $\mathcal{O}(N\log N)$ overhead in the worst case. This may challenge real-time performance requirements in dense traffic scenarios with numerous collaborative agents. In the future, we will explore more lightweight pre-screening methods to further reduce computational requirements of PASAC in dense collaborative networks, and expand the defensive capabilities against more sophisticated adversarial attacks by incorporating temporal consistency checks that leverage sequential data to identify anomalous behavior patterns.
7 Conclusion

### v3

@@ Abstract
Collaborative Perception (CP) has been shown to be a promising technique for multi-agent autonomous driving and multi-agent robotic systems, where multiple agents share their perception information to enhance the overall perception performance and expand the perception range. However, in CP, an ego agent needs to receive messages from its collaborators, which makes it vulnerable to attacks from malicious agents. To address this critical issue, we propose a unified, probability-agnostic, and adaptive framework, namely, CP-uniGuard, which is a tailored defense mechanism for CP deployed by each agent to accurately detect and eliminate malicious agents in its collaboration network. Our key idea is to enable CP to reach a consensus rather than a conflict against an ego agent’s perception results. Based on this idea, we first develop a probability-agnostic sample consensus (PASAC) method to effectively sample a subset of the collaborators and verify the consensus without prior probabilities of malicious agents. Furthermore, we define collaborative consistency loss (CCLoss) for object detection task and bird’s eye view (BEV) segmentation task to capture the discrepancy between an ego agent and its collaborators, which is used as a verification criterion for consensus. In addition, we propose online adaptive threshold via dual sliding windows to dynamically adjust the threshold for consensus verification and ensure the reliability of the systems in dynamic environments. Finally, we conduct extensive experiments and demonstrate the effectiveness of our framework. Code is available at https://github.com/CP-Security/CP-uniGuard.
@@ I Introduction
II Background and Related Work
II-A Collaborative Perception
@@ III Problem Setup
6:for $i=0$ to n1n-1 do
7:  $\mathbf{f}_{i,t}\leftarrow E_{\theta_{i}}(\mathbf{x}_{i,t})$ $\triangleright$ Extract features
@@
[…]
11:for $i=1$ to n1n-1 do
12:  $\mathbf{f}_{i,t}^{w}\leftarrow\mathcal{W}(\mathbf{f}_{i,t},\mathbf{T}_{i\to 0,t})$ $\triangleright$ Warp features
13:  $\mathbf{m}_{i,t}\leftarrow C_{\phi}(\mathbf{f}_{i,t}^{w})$ $\triangleright$ Compress features (optional)
14:  $\mathtt{Mess}_{t}\leftarrow\mathtt{Mess}_{t}\cup\{\mathbf{m}_{i,t}\}$ $\triangleright$ Add to message set
@@
[…]
@@
[…]
@@
III-A Collaborative Perception
As stated in Section II-A, CP can be categorized into three classes according to the stage of collaboration, including early fusion, intermediate fusion, and late fusion. In this section, we focus on the widely used intermediate fusion paradigm, which is the most efficient and effective fusion strategy.
@@ III-A Collaborative Perception
III-B Adversarial Threat Model in Collaborative Perception
@@
where $\boldsymbol{\delta}_{i,t}$ is the adversarial perturbation. $\mathcal{B}_{p}(\Delta)=\{x:\lVert x\rVert_{p}\leq\Delta\}$ is an $\ell_{p}$ ball to constrain the magnitude of the perturbation. Unless stated otherwise, we adopt the imperceptibility norm $p=\infty$ with $\Delta=0.1$ .
@@ IV CP-uniGuard Framework
14:  $\hat{\mathbf{y}}_{\mathrm{G}_{1},t}\leftarrow D_{\omega}(A_{\psi}(\mathbf{f}_{0,t},\{\mathbf{f}_{i,t}^{w}\}_{i\in\mathrm{G}_{1}}))$ $\triangleright$ Fusion
15:  $\hat{\mathbf{y}}_{\mathrm{G}_{2},t}\leftarrow D_{\omega}(A_{\psi}(\mathbf{f}_{0,t},\{\mathbf{f}_{i,t}^{w}\}_{i\in\mathrm{G}_{2}}))$ $\triangleright$ Fusion
@@
[…]
@@
[…]
@@
IV-A PASAC: Probability-Agnostic Sample Consensus
IV-A1 Our Method
@@
More specifically, the ego agent will generate the collaborative perception results $\mathbf{y}_{t}$ based on its observation and the received messages for feature fusion from the n1n-1 other agents. Firstly, the ego agent generates its perception results $\mathbf{y}_{0,t}$ based on its own observation. Then, it randomly splits the agents into two groups $\mathrm{G}_{1}$ and $\mathrm{G}_{2}$ of equal size (nearly equal when the number of agents is odd). After receiving all the messages, the ego agent fuses the features and generates the perception results $\mathbf{y}_{\mathrm{G}_{1},t}$ based on the messages from the first group and $\mathbf{y}_{\mathrm{G}_{2},t}$ based on the messages from the second group.
@@ IV-A2 Theoretical Analysis
In the worst-case scenario, we disregard the early stopping condition k=nmk=n-m, which is met when all benign agents are identified. We define the positive dyadic-logarithm $\bigl\lceil\log_{2}x\bigr\rceil_{+}$ as $\max(0,\lceil\log_{2}x\rceil)$.
@@
Let $L\triangleq\lceil\log_{2}n\rceil$. Fix a malicious agent $j\in\mathcal{M}$ and follow the unique root-to-leaf path that eventually isolates $\{j\}$. Starting from $|V|=n$, each split at most halves the size of the contaminated subset that contains $j$. After $L$ levels the subset size is <2m<2m; at most another $L$ splits are needed before every malicious singleton appears as a leaf. Hence, the depth of any contaminated path is bounded by $L$. There are $m$ such root-to-leaf paths (one per malicious agent). Each path contains at most $L$ contaminated nodes. Moreover, when a contaminated node $S$ is split, at most one additional contaminated sibling (the ”other half”) can be created. Therefore, the total number of contaminated internal nodes satisfies $|\textsc{Contam‐int}|\;<\;2mL$.
Every benign singleton $\{i\}$ with $i\notin\mathcal{M}$ is queried exactly once and produces a benign leaf. Hence there are precisely nmn-m such leaves. Adding the two contributions yields
@@
IV-B Collaborative Consistency Loss Verification
@@
IV-B1 Collaborative Consistency Loss for Object Detection
@@
IV-B2 Collaborative Consistency Loss for BEV Segmentation
@@
IV-C Online Adaptive Threshold via Dual Sliding Windows
@@
For any window $\mathcal{W}$ of size $n$, denote the empirical qq‑quantile by $\hat{q}_{q}(\mathcal{W})=\inf\bigl\{z:\tfrac{1}{n}\sum_{Z\in\mathcal{W}}\!\mathbf{1}\{Z\leq z\}\geq q\bigr\}.$ We use the upper $(1-\alpha)$‑quantile of benign scores and the lower $\beta$‑quantile of contaminated scores:
@@ V Experiments
TABLE I: Quantitative results on BEV segmentation task in V2X-Sim dataset. Upper-bound denotes collaborative perception with all benign agents. Lower-bound is individual perception.
@@
V-A Experimental Setup
Datasets and Evaluation Metrics. In our experiments, we leverage V2X-Sim [33] and DAIR-V2X [34] as our datasets, which are simulated and real-world datasets, respectively. In addition, to evaluate the performance of the object detection and BEV segmentation tasks, we adopt average precision (AP) with Intersection over Union (IoU) thresholds 0.5 and 0.7 (AP@0.5 and AP@0.7) for object detection task, and mean Intersection over Union (mIoU) for BEV segmentation task. We also use Verification Count to evaluate the performance of PASAC, which is the total number of times that malicious agents are checked.
Implementation Details. We leverage the backbone which is the same as [35] for the object detection task, and use U-Net [36] as the backbone for the BEV segmentation task. The fusion method for the both tasks is V2VNet [37]. Our experiment is deployed on a computer consisting of 2 Intel(R) Xeon(R) Silver 4410Y CPUs (2.0GHz), four NVIDIA RTX A5000 GPUs, and 512GB DDR4 RAM. As for the implementation of adversarial attacks, we employ three kinds of attacks: fast gradient sign method (FGSM) [38], Carlini & Wagner (C&W) [39], and the projected gradient descent (PGD) [40]. For each attack, we set the maximum perturbation $\delta_{\mathrm{max}}=0.1$, iterations steps $T=15$, and the step size $\gamma=0.01$.
@@ V-A Experimental Setup
V-B Evaluation and Analysis
V-B1 Evaluation of CP-uniGuard.
We evaluate the efficacy of our CP-uniGuard scheme against a variety of adversarial attacks. The outcomes of these evaluations are detailed in Table I, Table II, and Table III. As shown in Table I, in scenarios where the CP system lacks defensive mechanism, the mIoU across all three attack modalities significantly falls below the established lower bound, registering at 37.09%. This substantial degradation in performance underscores the effectiveness of the adversarial attacks implemented. Conversely, our CP-uniGuard framework can effectively defend against these attacks and achieve an mIoU that closely approaches the upper bound of 40.45%.
We also evaluate our approach on object detection (Table II). CP-uniGuard consistently achieves the best AP@0.5 and AP@0.7 under all attack scenarios. For example, against PGD, it obtains 80.4 (AP@0.5) and 78.3 (AP@0.7), outperforming ROBOSAC by 2.5 and 2.7 percentage points, respectively. For C&W attacks, the improvements over ROBOSAC are even larger (5.7 and 6.5 percentage points).
Finally, we evaluate our method in the real-world dataset, DAIR-V2X (Table III). As shown in the table, our method achieves the best performance among all compared defense methods, with AP@0.5 and AP@0.7 scores of 58.31% and 55.32%, respectively. This not only significantly outperforms both ROBOSAC and adversarial training, but also closes most of the gap to the upper bound where no attack is present. These results further validate the effectiveness and robustness of our approach on real-world datasets.
TABLE II: Quantitative results on object detection task in V2X-Sim dataset. The upper-bound represents collaborative perception with all benign agents, while the lower-bound corresponds to individual perception.
@@ V-B Evaluation and Analysis
TABLE III: Quantitative results on DAIR-V2X dataset.
Method AP@0.5 AP@0.7
No Defense (PGD Attack) 36.55 34.12
AdvTrain 50.21 47.45
ROBOSAC 42.19 40.52
CP-uniGuard (ours) 58.31 55.32
Upperbound 59.43 56.19
V-B2 Evaluation of PASAC.
@@
TABLE IV: Comparison results between ROBOSAC and PASAC. In this experiment, the attack ratios are known to the ROBOSAC method.
@@
TABLE V: Comparison of group checking latency.
Num Agent Method Latency per group checking (s)
5 ROBOSAC 1.62
PASAC 1.71
Linear 1.55
4 ROBOSAC 1.42
PASAC 1.47
Linear 1.34
3 ROBOSAC 1.02
PASAC 1.07
Linear 1.03
V-B3 Comparison Results in Sampling Efficiency.
We evaluate the sampling efficiency of PASAC compared to the previous state-of-the-art, ROBOSAC [5], following its experimental setup with known attack ratios. As shown in Table IV, PASAC consistently requires fewer verification steps than ROBOSAC across all attack ratios, and its results are notably more stable (e.g., much lower maximum verification count). We further compare PASAC, ROBOSAC, and Linear Sampling without prior knowledge of malicious agent ratios. As shown in Fig. 4, PASAC consistently requires fewer verification steps than the others, while ROBOSAC’s efficiency sharply drops when attack ratios are unknown. This highlights PASAC’s superior sampling efficiency and robustness.
Figure 5: Ablation study of the adaptive threshold adjustment dynamics for different initial threshold values and window sizes.
TABLE VI: Comparison between CCLoss and Mainstream Loss Functions
Detection Loss AP@0.5 AP@0.7
Smooth L1 [41] 49.64 48.79
IoU [42] 66.52 64.92
CCLoss (ours) 80.41 78.35
Segmentation Loss AP@IoU
Cross Entropy [43] 37.09
CCLoss (ours) 39.34
V-B4 Comparison Results in Computational Latency.
We measured the average latency per group checking [^1] This means the latency that verifies one collaborative group agents to identify the malicious agent. For example, with collaborative 5 agents, the latency of identifying the malicious agent from the 5 agents is used to denote this metric. for ROBOSAC, PASAC, and Linear checking, the results are shown in Table V. It demonstrates that the computational latency per group checking is roughly comparable across all methods. Since hardware constraints often necessitate processing agents in groups rather than all-at-once, PASAC’s ability to efficiently identify malicious agents through logarithmic search steps (fewer total checking times in large-scale networks) offers a better trade-off between efficiency and robustness compared to checking every agent individually.
V-B5 Comparison Results of CCLoss.
In addition, to further illustrate the effectiveness of our CCLoss, we conduct additional experiments to compare the CCLoss with mainstream loss functions, including Smooth L1 loss [41] and IoU loss [42] for object detection, and Cross Entropy loss [43] for segmentation. The results are shown in Table VI. We can see that CCLoss significantly outperforms the mainstream loss functions across both detection and segmentation tasks. These results demonstrate that our CCLoss not only provides a stronger verification criterion for consensus but also contributes directly to superior perception performance when compared to popular, widely-used alternatives.
TABLE VII: Sensitivity analysis of the initial threshold value of the online adaptive threshold.
Initial Threshold AP@0.5 AP@0.7
0.001 79.62 76.96
0.05 79.31 77.01
0.1 79.50 76.77
0.5 80.16 78.36
0.8 76.65 75.17
@@
Figure 6: Visualization of no defense and defensive CP-uniGuard and ROBOSAC results on V2X-Sim datasets.
V-B6 Study on the Online Adaptive Threshold Mechanism.
@@
In this section, we conducted an analysis on the online adaptive threshold mechanism. As shown in Fig. 5 (d), we plot the adaptive threshold value with different initial thresholds. We can observe that the threshold can quickly converge to a stable value, which is around 0.4. In addition, the threshold is not fixed, but dynamically adjusted based on the current situation to ensure the robustness of the system and a low false positive rate. Fig. 5 (a), (b), and (c) also indicate that the proposed mechanism consistently converges to a stable value regardless of initialization, and the window size controls the sensitivity of the threshold to the recent history of consistency scores, demonstrating robustness and fast adaptation.
In addition, Table VII presents results where we vary the initialization of the adaptive threshold across a wide range and report the corresponding performance. It shows that our framework maintains robustness to different initial values, as the average precision only varies within a small margin and the overall performance remains strong. This stable convergence to an effective value demonstrates that our adaptive adjustment mechanism mitigates the effect of this hyperparameter and further validates the robustness of our method with respect to threshold selection.
@@
TABLE VIII: Generalization Evaluation. We evaluate and compare the generalization capability of CP-uniGuard on two attacks, PGD and C&W, respectively.
@@
V-B7 Generalization Evaluation
We further evaluate the generalization capability of CP-uniGuard across different attack types, comparing it with feature-level defense methods such as CP-Guard+ [8]. For CP-Guard+, we train the model on the CP-GuardBench dataset [8] while excluding specific attacks from the training set to assess its generalization performance. Specifically, we exclude the PGD attack and C&W attack during training and subsequently evaluate the model’s performance against these two attacks, respectively. Results are presented in Table VIII, we can see that CP-uniGuard consistently achieves superior performance compared to CP-Guard+ under both PGD and C&W attacks which were unseen. This performance gap stems from CP-uniGuard’s fundamental design: rather than relying on training, CP-uniGuard leverages consensus verification, which provides inherent generalization to unseen attacks without requiring prior exposure during training.
V-B8 Qualitative Evaluation
As depicted in Fig. 6, we present the visualization results on the V2X-Sim dataset. Without CP-uniGuard, attackers can significantly disrupt collaborative perception, leading to a marked degradation in the performance of BEV segmentation tasks. However, our introduced CP-uniGuard framework can intelligently identify benign collaborators and eliminate malicious collaborators, thereby facilitating robust CP. In addition, compared with ROBOSAC which generates false positives and segmentation ambiguity, CP-uniGuard can achieve more accurate and stable results.
VI Limitations and Future Work
Despite the promising results of CP-uniGuard in defending against malicious agents in collaborative perception, our method still has some limitations and failure cases:
  1. 1.
    Adaptive Collusion in Blind Spots: If adaptive attackers coordinate to generate “consistent” false information (e.g., ghost objects) specifically located in the ego agent’s blind spot, they may evade the spatial consistency checks.
  2. 2.
    Adaptive Threshold Exploitation: While our online threshold $\varepsilon_{t}$ is dynamic, a sophisticated attacker with perfect knowledge of ego agent’s recent history could theoretically adapt their perturbation magnitude to stay marginally below the instantaneous rejection threshold.
  3. 3.
    Ego Sensor Failure: Since CP-uniGuard uses the ego agent as the root of trust, if an ego’s own sensors are severely compromised (e.g., by physical obstruction), the system may incorrectly flag benign collaborators as malicious due to the high discrepancy.
In our future work, we plan to address these limitations by enhancing the temporal and intrinsic robustness of the framework. To counter adaptive collusion in blind spots, we will incorporate multi-frame temporal consistency checks, leveraging historical trajectory data to distinguish valid dynamic agents from coordinated ghost objects that lack plausible motion continuity. To mitigate the risk of white-box threshold exploitation, we aim to introduce randomization mechanisms or long-term behavioral monitoring, enabling the detection of adversaries that persistently hover dangerously close to the rejection boundary. Finally, to address the single point of failure regarding ego sensor health, we will integrate intrinsic sensor health monitoring and cross-modal verification (e.g., LiDAR-Camera consistency), allowing the system to self-diagnose reliability issues and dynamically adjust confidence levels before penalizing collaborators.
VII Conclusion